# Accelerated Topological Pumping in Photonic Waveguides Based on Global Adiabatic Criteria

### v1

@@ Abstract
Adiabatic topological pumping promises robust transport of energy and information, but its speed is fundamentally limited by the instantaneous adiabatic condition, which demands prohibitively slow parameter variations. Here we develop a paradigm shift from instantaneous to global adiabaticity. We derive a global adiabatic criterion (GAC), which sets an absolute fidelity bound by controlling the root-mean-square value of nonadiabaticity factor. We further introduce a fluctuation-suppression acceleration criterion, which minimizes spatial inhomogeneity and allows us to safely increase the mean nonadiabaticity. Experimentally, we implement this principle in femtosecond-laser-written photonic Su-Schrieffer-Heeger waveguide arrays via scalable power-law coupling modulation. Our accelerated topological pumping achieves >0.95>0.95 fidelity over a fivefold reduced device length compared to the conventional scheme, exhibits the predicted linear scaling with the system size, and maintains robust performance across a >400>400 nm bandwidth. This principle of GAC provides a universal design rule for fast, compact, and robust adiabatic devices across quantum and classical topological platforms.
@@
Introduction—Topological photonics has emerged as a groundbreaking paradigm, offering robust energy transport immune to structural imperfections and disorder [1, 2, 3, 4]. This robustness arises from the topological protection of photonic edge states, which enables high-fidelity energy propagation even in the presence of defects, scattering, or fabrication variations [5, 6, 7, 8]. Notably, the well-known Su-Schrieffer-Heeger (SSH) model [9, 10] has become a cornerstone for studying phenomena such as topological pumping, where directed energy transfer occurs through adiabatic modulation of system parameters [11, 12, 13, 14, 15, 16, 17, 18, 19]. Recent experiments have demonstrated that topological pumping in SSH models of different platforms can achieve transports of energy on diverse physical carriers [20, 21, 22, 23, 24, 25, 26, 27].
@@
However, the practical implementation of such systems is fundamentally constrained by adiabatic protocols, which demand prohibitively large spatial footprints and conflict with the ongoing push toward device miniaturization [28, 29, 30, 31]. While various strategies have been explored to accelerate topological pumping, including including shortcut engineering techniques [32, 33, 34, 35, 36, 37], structure reforming [38, 20, 25, 21, 22, 23, 24, 39], and coupling tailoring [40, 41, 42, 43, 26, 27], they often optimize only local segments of the evolution, enhance experimental difficulties, or introduce sensitivity to parameter fluctuations. These limitations stem from the commonly considered instantaneous adiabatic condition, which imposes a pointwise constraint and overlooks the global statistical property of evolution. Consequently, it fails to provide a holistic design framework that can simultaneously ensure fast, high-fidelity transport across the entire pumping trajectory while maintaining scalability and robustness.
@@
In this work, we introduce a paradigm shift from instantaneous to global adiabaticity. We theoretically derive that high fidelity is governed not by suppressing nonadiabaticity at every point, but by controlling its root-mean-square (RMS) value. We develop two global accelerated adiabatic criteria (GAAC): a global adiabatic criterion (GAC) to enforce an absolute fidelity bound and a fluctuation-suppression acceleration criterion (FSAC) to minimize inhomogeneity penalties. This framework fundamentally rewrites the adiabatic constraint, permitting local rapid variation where the system is topologically robust, thereby enabling accelerated, linearly scalable pumping. We experimentally validate this paradigm in a femtosecond-laser-written photonic SSH lattice. By designing coupling profiles that satisfy our GAAC, we achieve high-fidelity topological pumping over a fivefold reduced device length compared to the conventional adiabatic scheme. In addition, we show that the accelerated topological pumping exhibits exceptional robustness across a >400>400 nm bandwidth. The core principle of GAAC is a fundamental strategy for accelerating any adiabatically slow, parameter-driven quantum or classical evolution. This includes, but is not limited to, quantum gate operations [44], Landau-Zener transitions [45], and other topological pumping protocols beyond one dimension [46].
@@
Theory framework—We start from the Schrödinger-like coupled-mode equation (CME) with respect to the spatial coordinate $z$ (consistent with our photonic waveguide platform) $\hat{H}(z)|\Psi(z)\rangle=i\partial_{z}|\Psi(z)\rangle$, where $|\Psi(z)\rangle$ is the system state vector. For a topological pumping system, we assume that the Hamiltonian $\hat{H}(z)$ possesses a non‑degenerate topological gap mode $|\mathrm{GAP}(z)\rangle$ with eigenenergy $E_{0}(z)$, separated from bulk modes $|\phi_{m}(z)\rangle$ ($m\neq 0$) by a finite energy gap $\Delta E_{m}(z)=|E_{m}(z)-E_{0}(z)|>0$. The conventional approach relies on the instantaneous adiabatic condition: $Q_{m}(z)\equiv\left|\frac{\langle\phi_{m}(z)|\partial_{z}|\mathrm{GAP}(z)\rangle}{\Delta E_{m}(z)}\right|\ll 1$, $\forall z\in[0,L]$ [30, 29]. This pointwise constraint enforces uniformly slow parameter variations, leading to prohibitively long device lengths.
@@
(i) Global adiabatic criterion [see Note 1 in Supplemental Material (SM[47] for detailed derivation]:
@@
GAC in Eq. (1) sets an absolute upper bound on the product of device length and the RMS value of nonadiabaticity factor, thereby guaranteeing the desired fidelity. FSAC in Eq. (2) eliminates the “penalty term” (fluctuation), in the cumulative nonadiabatic coupling by minimizing spatial inhomogeneity, allowing a larger mean driving strength ${\overline{Q}}_{\mathrm{non}}$ to be used without violating the adiabatic bound. When FSAC is satisfied, GAC simplifies to $L\,{\overline{Q}}_{\mathrm{non}}\,\Delta E_{\mathrm{min}}\leq\sqrt{\varepsilon_{c}}$. It reveals the central acceleration mechanism: for a given target fidelity and system gap, increasing the mean nonadiabaticity factor ${\overline{Q}}_{\mathrm{non}}$ (i.e., permitting faster parameter variations) directly shortens the required device length. Moreover, because the minimum gap $\Delta E_{\mathrm{min}}$ typically decreases with increasing system size $N$, our criteria inherently govern the scaling of $L$ with $N$.
To clarify our principle, in Fig. 1(a) we schematically show our acceleration strategy of topological pumping. $Q_{\mathrm{non}}(z)$ that obeys GAAC evolves more flatly with a relatively high mean value but smaller fluctuations, which can implement fast topological pumping with the target fidelity ${\mathcal{F}}_{c}$ by a critical evolution length $L_{c}$. By contrast, the conventional adiabatic design will spend a much longer length implementing ${\mathcal{F}}_{c}$ slowly, while failed when shortening length to $L_{c}$ due to enhanced nonadiabatic errors.
System and acceleration strategy—We execute exemplarily our investigation with the renowned dimerized SSH model [bottom panel in Fig. 1(b)]. In experiment, we fabricate the photonic waveguides (odd-numbered quantity is $N$) to construct the SSH model within a boroaluminosilicate glass by femtosecond-laser direct writing techniques [bottom panel in Fig. 1(c)] [48, 16, 49, 50]. This waveguide array with the length $L$ (maximum $75$ mm for our chips) is constrained to nearest-neighbor couplings, with staggered strengths $J_{1}$ (intracell) and $J_{2}$ (intercell) being spatially modulated in the $z$-direction.
@@
When the waveguide array is adiabatically modulated along the $z$ direction [29, 30, 14, 15, 25, 26, 27], the input light can be topologically pumped from one edge waveguide to the other following the spatial profiles of the zero-energy gap modes $|{\rm GAP}(z)\rangle={\mathcal{N}}\sum_{j=1}^{(N-1)/2}\beta^{j}(z)|2j+1\rangle$ [Note 2 in SM [47]], where $\beta(z)\equiv-J_{1}(z)/J_{2}(z)$ and $\mathcal{N}$} denotes the normalization. For the topologically nontrivial (trivial) phase under $J_{1}<J_{2}$ ($J_{1}>J_{2}$) (Note 3 in SM [47]), the gap mode is localized on the left (right), as shown by $|\Psi_{L(R)}\rangle$ in Fig. 1(b) (top panel). The topological pumping will be triggered when the waveguide array is adiabatically modulated from $J_{1}(z=0)\ll J_{2}(z=0)$ to $J_{1}(z=L)\gg J_{2}(z=L)$.
@@
In contrast, it is a quadratic function $J_{\rm max}L_{0.99}=1.22N^{2}+2.05N+5.15$ for the conventional pumping scheme based on $J_{1}(z)=0.5J_{\max}\left[1-\cos({\pi z}/{L})\right]$ and $J_{2}(z)=0.5J_{\max}\left[1+\cos({\pi z}/{L})\right]$ [14, 15, 42, 25, 27, 26]. This pair of coupling strengths yield the conventional adiabatic and nonadiabatic $Q_{\mathrm{non}}(z)$ in Fig. 1(a). This sharp contrast concerning the $L_{0.99}$-NN relationship underlines significant acceleration and scalibility superiority of our strategy over the conventional one.
To connect the $N$-dependent adaptive parameter $\alpha(N)$ with our GAAC, taking $N=13$ as an example (See Note 5 in SM [47] for sample fabrication), Fig. 2(c) displays comparison of ${\overline{Q}}_{\mathrm{non}}$ and $\sigma_{Q}^{2}$ between the conventional scheme and our strategy. On the one hand, our strategy exhibits significant suppression of ${\overline{Q}}_{\mathrm{non}}$ and $\sigma_{Q}^{2}$ compared to the conventional one, which is attributed to suppression of global nonadiabaticity. On the other hand, while ${\overline{Q}}_{\mathrm{non}}$ decreases with increasing $\alpha$ (with a decreasing rate), $\sigma_{Q}^{2}$ displays an apparent basin around $\alpha=3$ [there is a slight deviation from $\alpha(13)=3$, because $\alpha(N)$ is determined quantitatively by $L_{0.99}$ instead of $\sigma_{Q}^{2}$]. Referring to Fig. 2(d), a pumping fidelity peak exactly indicates optimal values around $\alpha=3$ that can provide a shorter device length for implementing the same fidelity. The critical length is $L_{0.95}=15$ mm for implementing ${\mathcal{F}}(L)>0.95$, $\forall L>L_{0.95}$ based on our GAAC with $\alpha(13)=3$, while it is $75$ mm for the conventional scheme, indicating a fivefold reduction in device footprint.
Existing schemes [40, 41, 40, 42, 43, 26, 27] for accelerating topological pumping may conform our GAAC to a certain degree. However, these schemes either do not fully adhere to both criteria simultaneously or lack a holistic framework that guarantees performance (speed and robustness) and scalability ($N$-dependent parameters). Specially, we consider comparison with recently reported fast topological pumping schemes (Note 6 in SM [47]), such as the coherent adiabatic passage with $J_{\rm max}L|_{N=31}=400$ for ${\mathcal{F}}>0.99$ [38, 20], the quantum metric with $J_{\rm max}L|_{N=8}=44.6$ for ${\mathcal{F}}>0.9$ [25], and the adiabatic infimum with $J_{\rm max}L|_{N=8}=56.8$ for ${\mathcal{F}}>0.9$ [27]. Our acceleration strategy for implementing ${\mathcal{F}}>0.99$ requires even more compact footprints $J_{\rm max}L|_{N=31}=70$ and $J_{\rm max}L|_{N=9}=40$. This remarkable contrast clearly demonstrates the superiority of our strategy in terms of topological pumping speed.
@@
Experimental demonstration—To validate the acceleration performance, we experimentally fabricated three sets of 13-waveguide samples with lengths $L=35$, $55$, and $75$ mm, respectively (Note 5 in SM [47]). Each set contains two samples corresponding to our acceleration strategy and the conventional scheme. An 808 nm wavelength laser source (CNI, MDL-III-808L) was coupled into the first waveguide of each sample as the input port. The resulting light intensity patterns were captured at the output side of each sample using a charge-coupled device (SP928, Ophir). Figures 3(a)-(c) (right panels) present the measured light intensity photographs of the end facets for the six samples. The measured output intensity distributions across the 13 waveguides in each sample are in good agreement with their respective theoretical values. According to both measured and predicted results, all the three samples for our acceleration strategy show almost fully localized light on the 13th waveguide representing the other edge of the waveguide array. On the contrary, for the conventional adiabatic pumping samples, the light intensity on the last waveguide is enhanced with increased sample length. Overall, significant light diffusion is observed in the bulk of the waveguide array of these conventional adiabatic samples, but our accelerated topological pumping sample of the same length exhibits strong light localization on the last waveguide. These results clearly demonstrate that our GAAC enables fast topological pumping for photonic transport with remarkably reduced device length and minimal bulk diffusion.
@@
Conclusion—We have theoretically formulated and experimentally verified an efficient strategy for accelerating adiabatic topological pumping by developing GAAC, transcending the limitations of conventional instantaneous conditions. By deriving that fidelity is governed by the spatial mean ${\overline{Q}}_{\rm non}$ and variance $\sigma^{2}_{Q}$ of the nonadiabatic factor, we enable rapid parameter modulation where the topological gap is large, while maintaining adiabaticity globally. Implemented in femtosecond-laser-written photonic lattices via scalable power-law couplings, this approach achieves high-fidelity edge-to-edge transport in device lengths much shorter than the conventional pumping scheme, exhibits the predicted linear scalability ($L\propto N$), and maintains robustness over a broad optical bandwidth (>400>400 nm). Moreover, as detailed by Note 6 in SM [47], our approach also surpasses other recently reported acceleration strategies [25, 27, 20] in key metrics.
As our extended experiment detailed by Note 8 in SM [47], the same framework readily extends to functional devices, such as topological beam splitters. The specific power-law coupling functions presented here serve as a representative example that satisfies our GAAC. In principle, any modulation profile conforming to these criteria is applicable, underscoring the universality of the framework. Moreover, our analytical GAAC can be integrated with numerical optimization methods (e.g., the recently reported adiabaticity control limit by gradient descent method [51] or time-optimal state transfer by shooting method [52]) to systematically navigate the full design space. This synergy promises to further unleash the potential for accelerated and robust topological pumping, as well as other adiabatic-evolution tasks, across diverse physical platforms from ultracold atoms [53, 44] to superconducting circuits [21, 54].
Acknowledgments—This work was supported in part by the National Natural Science Foundation of China under Grant Nos. 82003284 and 62375103, the National Key Research and Development Program of China under Grant Nos. 2024YFB4608100, 2021YFF0502700 and 2023YFB3208600. J.-L.W. acknowledges support from the National Natural Science Foundation of China (62571494,12304407). X.N. acknowledges support from the National Natural Science Foundation of China (12474396), and Key Project of Hunan Natural Science Foundation (2025JJ30001).

### v2

@@ Abstract
Adiabatic topological pumping enables robust transport of energy and information, yet its operational speed is fundamentally constrained by the instantaneous adiabatic condition, which necessitates prohibitively slow parameter variations. Here, we propose a paradigm shift from instantaneous to global adiabaticity. We derive a global adiabatic criterion (GAC) that establishes an absolute fidelity bound by controlling the root-mean-square nonadiabaticity. Building on this framework, we introduce a fluctuation-suppression acceleration criterion to minimize spatial inhomogeneity, allowing for a safe increase in mean nonadiabaticity without compromising fidelity. We experimentally demonstrate this principle in femtosecond-laser-written photonic Su-Schrieffer-Heeger waveguide arrays via scalable power-law coupling modulation. Our accelerated topological pumping achieves a fidelity of >0.95>0.95 with a fivefold reduction in device length compared to conventional schemes, exhibits the predicted linear scaling with system size, and maintains robust performance across a bandwidth exceeding $400$ nm. This GAC framework provides a universal design rule for fast, compact, and robust adiabatic devices across both quantum and classical topological platforms.
@@
Introduction—Topological photonics has emerged as a groundbreaking paradigm, offering robust energy transport immune to structural imperfections and disorder [1, 2, 3, 4]. This robustness arises from the topological protection of photonic edge states, which enables high-efficiency energy propagation even in the presence of defects, scattering, or fabrication variations [5, 6, 7, 8]. Notably, the well-known Su-Schrieffer-Heeger (SSH) model [9, 10] has become a cornerstone for studying phenomena such as topological pumping, in which directed energy transfer is achieved through adiabatic modulation of system parameters [11, 12, 13, 14, 15, 16, 17, 18, 19]. Recent experiments have demonstrated that SSH-based topological pumping across various platforms can facilitate energy transport using diverse physical carriers [20, 21, 22, 23, 24, 25, 26, 27].
@@
However, the practical implementation of such systems is fundamentally constrained by adiabatic protocols, which demand prohibitively large spatial footprints and conflict with the ongoing push toward device miniaturization [28, 29, 30, 31]. While various strategies have been explored to accelerate topological pumping, including shortcut engineering techniques [32, 33, 34, 35, 36, 37], structure reforming [38, 20, 25, 21, 22, 23, 24, 39], and coupling tailoring [40, 41, 42, 43, 26, 27], they often optimize only local segments of the evolution, exacerbate experimental complexities, or increase susceptibility to parameter fluctuations. These limitations stem from the commonly adopted instantaneous adiabatic condition, which imposes a pointwise constraint while overlooking the global statistical properties of evolution. Consequently, existing approaches fail to provide a holistic design framework capable of simultaneously ensuring fast, high-fidelity transport across the entire pumping trajectory while maintaining scalability and robustness.
@@
In this work, we introduce a paradigm shift from instantaneous to global adiabaticity. We theoretically show that high fidelity is governed not by suppressing nonadiabaticity at every point, but by controlling its root-mean-square (RMS) value over the entire evolution. Based on this insight, we develop two global accelerated adiabatic criteria (GAAC): a global adiabatic criterion (GAC) that enforces an absolute fidelity bound, and a fluctuation-suppression acceleration criterion (FSAC) that minimizes penalties arising from spatial inhomogeneity. This framework fundamentally reformulates the conventional adiabatic constraint, permitting local rapid parameter variations in topologically protected region, thereby enabling accelerated, linearly scalable pumping. We experimentally validate this paradigm in a femtosecond-laser-written photonic SSH lattice. By designing coupling profiles that satisfy the GAAC, we achieve high-fidelity topological pumping with a fivefold reduction in device length compared to conventional adiabatic scheme. In addition, the accelerated topological pumping exhibits exceptional robustness over a bandwidth exceeding $400$ nm. Beyond topological pumping, the GAAC principle provide a general strategy for accelerating any adiabatically driven quantum and classical evolutions, including quantum gate operations [44, 45], Landau-Zener transitions [46, 47], and higher-dimensional topological pumping protocols [48, 49].
@@
Theory framework—We start from the Schrödinger-like coupled-mode equation (CME) defined with respect to the spatial coordinate $z$ (consistent with our photonic waveguide platform) $\hat{H}(z)|\Psi(z)\rangle=i\partial_{z}|\Psi(z)\rangle$, where $|\Psi(z)\rangle$ is the system state vector. For a topological pumping system, we assume that the Hamiltonian $\hat{H}(z)$ possesses a non‑degenerate topological gap mode $|\mathrm{GAP}(z)\rangle$ with eigenenergy $E_{0}(z)$, separated from the $m$th bulk mode $|\phi_{m}(z)\rangle$ ($m\neq 0$) by a finite energy gap $\Delta E_{m}(z)=|E_{m}(z)-E_{0}(z)|>0$. The conventional approach relies on the instantaneous adiabatic condition: $Q_{m}(z)\equiv\left|\frac{\langle\phi_{m}(z)|\partial_{z}|\mathrm{GAP}(z)\rangle}{\Delta E_{m}(z)}\right|\ll 1$, $\forall z\in[0,L]$ [30, 29]. This pointwise constraint enforces uniformly slow parameter variations, leading to prohibitively long device lengths.
@@
(i) Global adiabatic criterion [see Note 1 in Supplemental Material (SM) [50] for detailed derivation]:
@@
GAC in Eq. (1) sets an absolute upper bound on the product of device length and the RMS value of nonadiabaticity factor, thereby guaranteeing the desired fidelity. FSAC in Eq. (2) eliminates the “penalty term” arising from fluctuation, in the cumulative nonadiabatic coupling by minimizing spatial inhomogeneity, thereby permitting a larger mean driving strength ${\overline{Q}}_{\mathrm{non}}$ without violating the adiabatic bound. When FSAC is satisfied, GAC simplifies to $L\,{\overline{Q}}_{\mathrm{non}}\,\Delta E_{\mathrm{min}}\leq\sqrt{\varepsilon_{c}}$. It reveals the central acceleration mechanism: for a given target fidelity and system gap, increasing the mean nonadiabaticity factor ${\overline{Q}}_{\mathrm{non}}$ (i.e., permitting faster parameter variations) directly shortens the required device length. Moreover, because the minimum gap $\Delta E_{\mathrm{min}}$ typically decreases with increasing system size $N$, our criteria inherently govern the scaling of $L$ with $N$.
To illustrate our principle, Fig. 1(a) schematically depicts the proposed acceleration strategy for topological pumping. The nonadiabaticity $Q_{\mathrm{non}}(z)$ satisfying GAAC evolves more uniformly, maintaining a relatively large mean value while exhibiting reduced fluctuations. This profile enables fast topological pumping to achieve target fidelity ${\mathcal{F}}_{c}$ within a critical evolution length $L_{c}$. In contrast, conventional adiabatic designs require a substantially longer device length to reach the same fidelity, and fail to do so when compressed to $L_{c}$ due to enhanced nonadiabatic errors.
System and acceleration strategy—We execute exemplarily our investigation with the renowned dimerized SSH model [bottom panel in Fig. 1(b)]. In experiment, we fabricate the photonic waveguides (odd-numbered quantity is $N$) to construct the SSH model within a boroaluminosilicate glass by femtosecond-laser direct writing techniques [Fig. 1(c)] [51, 16, 52, 53]. This waveguide array with the length $L$ (maximum $75$ mm for our chips) is constrained to nearest-neighbor couplings, with staggered strengths $J_{1}$ (intracell) and $J_{2}$ (intercell) being spatially modulated in the $z$-direction.
@@
When the waveguide array is adiabatically modulated along the $z$ direction [29, 30, 14, 15, 25, 26, 27], the input light can be topologically pumped from one edge waveguide to the other following the spatial profiles of the zero-energy gap modes $|{\rm GAP}(z)\rangle={\mathcal{N}}\sum_{j=1}^{(N-1)/2}\beta^{j}(z)|2j+1\rangle$ (Note 2 in SM [50]), where $\beta(z)\equiv-J_{1}(z)/J_{2}(z)$ and $\mathcal{N}$} denotes the normalization. For the topologically nontrivial (trivial) phase under $J_{1}<J_{2}$ ($J_{1}>J_{2}$) (Note 3 in SM [50]), the gap mode is localized on the left (right), as shown by $|\Psi_{L(R)}\rangle$ in Fig. 1(b) (top panel). The topological pumping will be triggered when the waveguide array is adiabatically modulated from $J_{1}(z=0)\ll J_{2}(z=0)$ to $J_{1}(z=L)\gg J_{2}(z=L)$.
@@
In contrast, it is a quadratic function $J_{\rm max}L_{0.99}=1.22N^{2}+2.05N+5.15$ for the conventional pumping scheme based on $J_{1}(z)=0.5J_{\max}\left[1-\cos({\pi z}/{L})\right]$ and $J_{2}(z)=0.5J_{\max}\left[1+\cos({\pi z}/{L})\right]$ [14, 15, 42, 25, 27, 26]. This pair of coupling strengths yields the conventional adiabatic and nonadiabatic profiles of $Q_{\mathrm{non}}(z)$ shown in Fig. 1(a). This sharp contrast concerning the $L_{0.99}$-NN relationship underlines significant acceleration and scalibility superiority of our strategy over the conventional one. Distinct enlargement of $\Delta E_{\rm min}$ [Fig. 2(b)] in our acceleration strategy is an important factor resulting in a much shorter $L_{0.99}$ than the conventional scheme, which can further decrease the upper bound of evolution length according to GAC in Eq. (1).
To connect the $N$-dependent adaptive parameter $\alpha(N)$ with our GAAC, we take $N=13$ as an example (See Note 5 in SM [50] for sample fabrication). Figure 2(c) compares ${\overline{Q}}_{\mathrm{non}}$ and $\sigma_{Q}^{2}$ between the conventional scheme and our strategy. On the one hand, our strategy achieves significant suppression of both ${\overline{Q}}_{\mathrm{non}}$ and $\sigma_{Q}^{2}$ compared to the conventional approach, which is attributed to overall reduction of global nonadiabaticity. On the other hand, while ${\overline{Q}}_{\mathrm{non}}$ decreases with increasing $\alpha$ (with a diminishing rate of decrease), $\sigma_{Q}^{2}$ exhibits a clear minimum around $\alpha=3$. Referring to Fig. 2(d), the peak in pumping fidelity identifies an optimal value around $\alpha=3$, which enables a significantly shorter device length to achieve the same target fidelity. Under the GAAC design with $\alpha(13)=3$, the critical length required to achieve ${\mathcal{F}}(L)>0.95$ is $L_{0.95}=15$ $m$, $\forall L>L_{0.95}$. In contrast, the conventional scheme requires a length of 75 mm to reach the same fidelity, indicating a fivefold reduction via our scheme in device footprint.
Existing schemes [40, 41, 40, 42, 43, 26, 27] for accelerating topological pumping may partially align with our GAAC. However, these schemes either do not fully satisfy both criteria simultaneously or lack a comprehensive framework that guarantees performance (speed and robustness) and scalability (via $N$-dependent parameters). Specially, we focus on comparison with recently reported fast topological pumping schemes (Note 6 in SM [50]), such as the coherent adiabatic passage with $J_{\rm max}L|_{N=31}=400$ for ${\mathcal{F}}>0.99$ [38, 20], the quantum metric with $J_{\rm max}L|_{N=8}=44.6$ for ${\mathcal{F}}>0.9$ [25], and the adiabatic infimum with $J_{\rm max}L|_{N=8}=56.8$ for ${\mathcal{F}}>0.9$ [27]. Our acceleration strategy for implementing ${\mathcal{F}}>0.99$ requires even more compact footprints $J_{\rm max}L|_{N=31}=70$ and $J_{\rm max}L|_{N=9}=40$. This remarkable contrast clearly demonstrates the superiority of our strategy in terms of topological pumping speed.
@@
Experimental demonstration—To validate the acceleration performance, we experimentally fabricated three sets of 13-waveguide samples with lengths $L=35$, $55$, and $75$ mm, respectively (Note 5 in SM [50]). Each set contains two samples corresponding to our acceleration strategy and the conventional scheme. An 808 nm wavelength laser source (CNI, MDL-III-808L) was coupled into the first waveguide of each sample as the input port. The resulting light intensity patterns were captured at the output side of each sample using a charge-coupled device (SP928, Ophir). Figures 3(a)-(c) (right panels) present the measured light intensity photographs of the end facets for the six samples. The measured output intensity distributions across the 13 waveguides in each sample are in excellent agreement with their respective theoretical values. According to both measured and predicted results, all the three samples for our acceleration strategy show almost fully localized light on the 13th waveguide representing the other edge of the waveguide array. On the contrary, for the conventional adiabatic pumping samples, the light intensity on the last waveguide is enhanced with increased sample length. Overall, significant light diffusion is observed in the bulk of the waveguide array of these conventional adiabatic samples, but our accelerated topological pumping sample of the same length exhibits strong light localization on the last waveguide. These results clearly demonstrate that our GAAC enables fast topological pumping for photonic transport with remarkably reduced device length and minimal bulk diffusion.
@@
Conclusion—We have theoretically formulated and experimentally verified a high-efficiency framework for accelerating adiabatic topological pumping via GAAC, transcending the limitations of conventional instantaneous adiabatic conditions. By establishing that fidelity is fundamentally governed by the spatial mean ${\overline{Q}}_{\rm non}$ and variance $\sigma^{2}_{Q}$ of the nonadiabatic factor, we have facilitated rapid parameter modulation where the topological gap is large while preserving global adiabaticity. Using femtosecond-laser-written photonic lattices with scalable power-law couplings modulations, we have demonstrated high-fidelity, edge-to-edge transport with device footprint significantly shorter than those required by conventional pumping schemes. Our results exhibit the predicted linear scalability ($L\propto N$), and maintains exceptional robustness over a broad optical bandwidth (>400>400 nm). Moreover, as detailed by Note 6 in SM [50], our approach also surpasses other recently reported acceleration strategies [25, 27, 20] in key metrics.
The versatility of GAAC framework is further evidenced by its ready extension to functional devices, such as topological beam splitters (see SM, Note 8 [50]). While we have employed power-law coupling modulations as a primary realization, the GAAC is agnostic to the specific functional form, highlighting its broad applicability. This analytical GAAC can be synergistically integrated with numerical optimization methods (e.g., the gradient-descent-based adiabaticity control limit [54], time-optimal shooting method [55], or optimal-pulse genetic algorithm [56]) to systematically navigate the full design space. Such a synergy of analytical bounds and numerical refinement may provide a robust pathway for maximizing the speed and fidelity of topological transport, as well as broader adiabatic protocols, across diverse physical platforms from ultracold atoms [57, 58, 45] to superconducting circuits [21, 59, 60].
Acknowledgments—This work was supported in part by the National Natural Science Foundation of China under Grant Nos. 82003284 and 62375103, the National Key Research and Development Program of China under Grant Nos. 2024YFB4608100, 2021YFF0502700 and 2023YFB3208600. J.-L.W. acknowledges support from the National Natural Science Foundation of China (62571494,12304407). X.N. acknowledges support from the National Natural Science Foundation of China (12474396), and Key Project of Hunan Natural Science Foundation (2025JJ30001). Y.-K.S. acknowledges support from the National Natural Science Foundation of China (623B2042). K.-H.X. thanks X. Yang and W.-Z. Fan for their assistance with the sample fabrication and measurement.